# GL1-02-06-24-2-supplement

### v2

@@ 7.2 Estimation With Underreported Outcomes
$-\frac{1}{n_{11}}\sum_{i=1}^{n_{11}}\mathbb{I}\left[G_{i}=1,T_{i}=1\right]\mathbb{I}\left[\mathbf{X}_{i}\in\mathcal{X}_{11}\right]\left(F_{-\varepsilon}\left(\frac{\overline{\mathbf{X}}_{i}^{\prime}\widehat{\overline{\boldsymbol{\eta}}}^{(0)}_{11}-\kappa_{j+1}}{\widehat{\lambda}^{(0)}_{11}}\right)-F_{-\varepsilon}\left(\frac{\overline{\mathbf{X}}_{i}^{\prime}\widehat{\overline{\boldsymbol{\eta}}}^{(0)}_{11}-\kappa_{j}}{\widehat{\lambda}^{(0)}_{11}}\right)\right).$
The asymptotic distribution of these DTT variants can be derived similarly to the approach in Subsection 7.1, with standard errors constructed accordingly.
Remark 3: Test statistics for the hypotheses $H_{0}^{C}$ vs. $H_{A}^{C}$ or $H_{0}^{R}$ vs. $H_{A}^{R}$, as discussed in Remark 1, may be readily constructed using standard Cramer-von Mises measures of distance. Specifically, for testing $H_{0}^{C}$ vs. $H_{A}^{C}$, the following statistic can be employed for some $(g.t)\neq(1,1)$:
$\sum_{i=1}^{n}\sum_{j=1}^{2}\mathbb{I}\left[G_{i}=g,T_{i}=t\right]\mathbb{I}\left[(\mathbf{X}_{i},\mathbf{Z}_{i})\in\mathcal{X}_{gt}\times\mathcal{Z}_{gt}\right]$
×$\times\left[\overline{C}\left(F_{-{\varepsilon}}\left(\frac{\overline{\mathbf{X}}_{i}^{\prime}\widehat{\overline{\boldsymbol{\eta}}}_{gt}^{(0)}-\kappa_{j}}{\widehat{\lambda}_{gt}^{(0)}}\right),F_{-{\nu}}\left(\frac{\overline{\mathbf{Z}}_{i}^{\prime}\widehat{\overline{\boldsymbol{\pi}}}_{gt}^{(0)}-\iota_{j}}{\widehat{\zeta}_{gt}^{(0)}}\right);\widehat{\rho}_{gt}^{(0)}\right)-\widehat{M}\left(j;\mathbf{X}_{i}^{\prime}\widehat{\widetilde{\boldsymbol{\eta}}}_{gt},\mathbf{Z}_{i}^{\prime}\widehat{\widetilde{\boldsymbol{\pi}}}_{gt}\right)\right]^{2},$
where $\left(\widehat{\overline{\boldsymbol{\eta}}}_{gt}^{(0)^{\prime}},\widehat{\lambda}_{gt}^{(0)},\widehat{\overline{\boldsymbol{\pi}}}_{gt}^{(0)^{\prime}},\widehat{\zeta}^{(0)}_{gt},\widehat{\rho}_{gt}^{(0)}\right)$ are ML estimators discussed before, $\left(\widehat{\widetilde{\boldsymbol{\eta}}}_{gt}^{\prime},\widehat{\widetilde{\boldsymbol{\pi}}}_{gt}^{\prime}\right)$ denote multi-index estimators (Donkers and Schafgans, 2008; Ahn et al., 2018), and $\widehat{M}(j;\cdot,\cdot)$ is a suitable nonparametric estimator. A similar approach can be formulated for testing $H_{0}^{R}$ vs. $H_{A}^{R}$. The limiting distributions of both tests, while beyond the scope of the paper, are standard, and suitable resampling procedures can be used to construct asymptotically valid critical values.
8 Simulation
In this section, we present Monte Carlo simulation evidence to investigate the finite sample behavior of the ML estimator outlined in Subsection 7.2. Our focus is on assessing the finite sample bias and variance of the ML estimator, specifically in relation to the model parameters for marginal consumption and reporting distributions, as well as the dependence parameter in the copula function. The analysis spans various exclusion restriction scenarios, comparing cases with equal index and error variances to those where the index variance is less than the error variance.
The design of our observed outcome equation for simulation aligns with that in (1), but streamlined by not incorporating the treatment effects framework and simplifying the notation for the potential outcomes and the $gt$ subscript. Specifically, we model the consumption and reporting decisions as:
$C=\sum_{j=0}^{2}\mathbb{I}\left[\beta_{0}+\beta_{1}X+\beta_{2}W-\lambda\varepsilon>\kappa_{j}\right],$
$R=\sum_{j=0}^{2}\mathbb{I}\left[\pi_{0}+\pi_{1}Z+\pi_{2}W-\zeta\nu>\iota_{j}\right].$
Here, $W\sim\mathcal{N}(0,1)$ represents a common regressor, while $X$ and $Z$ are the “excluded” variables following a uniform distribution $\mathcal{U}\left(-\sqrt{3},\sqrt{3}\right)$ (detailed in Table 1). The joint distribution $(\varepsilon,\nu)^{\prime}$ is modeled using a Frank copula, denoted as $\overline{C}\left(F_{\varepsilon},F_{\nu};\rho\right)$, with the marginals being standard normal distributions. The Frank copula, representative of symmetric copulas, is defined as:
$\overline{C}(v_{1},v_{2};\rho)=-\frac{1}{\rho}\ln\left(1+\frac{(\exp(-\rho v_{1})-1)(\exp(-\rho v_{2})-1)}{\exp(-\rho)-1}\right)$
for $\rho\neq 0$. When $\rho=0$, it defaults to the independence copula.
For the simulation purposes, the dependence parameter $\rho$ is set such that the corresponding Spearman’s $\rho$ is 0.50.5-0.5- 0.5.111111Results for cases where the Spearman’s $\rho$ equals $-0.25$, yielding qualitatively similar outcomes, are available in Subsection LABEL:ssec:frank of the supplementary material. This approach is intended to facilitate comparability between dependence parameters across copulas, each of which can have distinct interpretations (see Han and Lee, 2019). Additionally, in Section LABEL:sec:simulations of the supplementary material, we extend our analysis to include simulations with the Clayton Copula, known for its asymmetric properties. These simulations yield findings that are qualitatively similar to those obtained using the Frank copula, thus highlighting the robustness of our study. We consider sample sizes $n=\{2000,5000,10000\}$, with each size undergoing 2000 Monte Carlo replications. To assess finite sample bias and variance, we focus on the Mean Bias (MB) and the Root Mean Square Error (RMSE).
Table 1 provides an overview of the various designs examined in our Monte Carlo study. We selected distributions and parameter values to ensure that the (unconditional) variances of the indices and the unobservables are equal to each other, with the exceptions of Cases 1-b and 2-b.
Table 1: Simulation Design Cases
Case No. of Exclusions Specification
0 No Exclusion $C=\sum_{j=1}^{2}\mathbb{I}\left[-1+2\cdot W-2\cdot\varepsilon>\kappa_{j}\right]$
$R=\sum_{j=1}^{2}\mathbb{I}\left[1.5-2\cdot W-2\cdot\nu>\iota_{j}\right]$
1-a 1 Exclusion $C=\sum_{j=1}^{2}\mathbb{I}\left[-1+2\cdot W-2\cdot\varepsilon>\kappa_{j}\right]$
$R=\sum_{j=1}^{2}\mathbb{I}\left[1.5-\sqrt{2}\cdot W-\sqrt{2}\cdot Z-2\cdot\nu>\iota_{j}\right]$
1-b 1 Exclusion $C=\sum_{j=1}^{2}\mathbb{I}\left[-1+\sqrt{2}\cdot W-2\cdot\varepsilon>\kappa_{j}\right]$
$R=\sum_{j=1}^{2}\mathbb{I}\left[1.5-W-Z-2\cdot\nu>\iota_{j}\right]$
2-a 2 Exclusions $C=\sum_{j=1}^{2}\mathbb{I}\left[-1+\sqrt{2}\cdot W+\sqrt{2}\cdot X-2\cdot\varepsilon>\kappa_{j}\right]$
$R=\sum_{j=1}^{2}\mathbb{I}\left[1.5-\sqrt{2}\cdot W-\sqrt{2}\cdot Z-2\cdot\nu>\iota_{j}\right]$
2-b 2 Exclusions $C=\sum_{j=1}^{2}\mathbb{I}\left[-1+W+X-2\cdot\varepsilon>\kappa_{j}\right]$
$R=\sum_{j=1}^{2}\mathbb{I}\left[1.5-W-Z-2\cdot\nu>\iota_{j}\right]$
Case 0 represents a scenario without exclusion restrictions, where identification is expected to fail. In contrast, Cases 1-a and 1-b focus on designs with a single exclusion restriction, differing in their (unconditional) index variances. Case 1-a considers equal (unconditional) index and error variances, whereas Case 1-b explores a setting with a smaller index variance relative to that of the error. Extending this framework, Cases 2-a and 2-b, akin to Cases 1-a and 1-b, respectively, incorporate two distinct exclusion restrictions in each equation.
Table 2: Frank Copula - Mean Bias
Case 0
Param. obs = 2000, $\rho$ = -0.5 obs = 5000, $\rho$ = -0.5 obs = 10000, $\rho$ = -0.5
$\beta_{0}$ $45.387$$10.271$$0.289$
$\beta_{1}$ $605.906$ $92.359$ $16.638$
$\pi_{0}$ $62.434$$12.121$$1.343$
$\pi_{1}$ $651.649$ $141.740$ $2.395$
$\lambda$ $719.680$ $79.700$ $16.190$
$\zeta$ $715.016$ $138.753$ $0.252$
$\rho$ $0.393$ $0.300$ $0.218$
Case 1-a
Param. obs = 2000, $\rho$ = -0.5 obs = 5000, $\rho$ = -0.5 obs = 10000, $\rho$ = -0.5
$\beta_{0}$$9.946$ $0.011$ $0.007$
$\beta_{1}$ $124.558$ $0.110$ $0.046$
$\pi_{0}$ $100.666$$0.229$$0.092$
$\pi_{1}$$77.552$ $0.110$ $0.040$
$\pi_{2}$$74.046$ $0.020$$0.002$
$\lambda$ $115.175$ $0.076$ $0.031$
$\zeta$ $91.241$$0.021$ $0.006$
$\rho$ $0.253$ $0.144$ $0.066$
Case 1-b
Param. obs = 2000, $\rho$ = -0.5 obs = 5000, $\rho$ = -0.5 obs = 10000, $\rho$ = -0.5
$\beta_{0}$$245.968$ $4.973$$0.011$
$\beta_{1}$ $1,902.489$ $15.519$ $0.128$
$\pi_{0}$ $252.689$$0.445$$0.341$
$\pi_{1}$$154.198$ $0.178$ $0.138$
$\pi_{2}$$149.381$ $0.034$ $0.040$
$\lambda$ $2,350.347$ $20.812$ $0.160$
$\zeta$ $270.014$$0.080$$0.078$
$\rho$ $0.364$ $0.273$ $0.190$
Case 2-a
Param. obs = 2000, $\rho$ = -0.5 obs = 5000, $\rho$ = -0.5 obs = 10000, $\rho$ = -0.5
$\beta_{0}$$0.035$$0.019$$0.005$
$\beta_{1}$ $0.114$ $0.046$ $0.016$
$\beta_{2}$ $0.118$ $0.046$ $0.017$
$\pi_{0}$ $13.961$ $0.043$ $0.028$
$\pi_{1}$$9.951$$0.040$$0.023$
$\pi_{2}$$11.006$$0.038$$0.024$
$\lambda$ $0.133$ $0.052$ $0.020$
$\zeta$ $13.324$ $0.043$ $0.029$
$\rho$ $0.026$ $0.007$ $0.001$
Case 2-b
Param. obs = 2000, $\rho$ = -0.5 obs = 5000, $\rho$ = -0.5 obs = 10000, $\rho$ = -0.5
$\beta_{0}$$280.768$$3.627$$0.014$
$\beta_{1}$ $603.093$ $9.119$ $0.055$
$\beta_{2}$ $603.862$ $8.056$ $0.050$
$\pi_{0}$ $238.591$ $0.152$ $0.063$
$\pi_{1}$$139.640$$0.088$$0.038$
$\pi_{2}$$147.751$$0.096$$0.034$
$\lambda$ $1,122.068$ $16.560$ $0.091$
$\zeta$ $285.927$ $0.167$ $0.071$
$\rho$ $0.063$ $0.017$ $0.004$
Notes: See Table 1 for the specific design of each case. All results are produced using $2000$ Monte Carlo replications.
Table 3: Frank Copula - RMSE
Case 0
Param. obs = 2000, $\rho$ = -0.5 obs = 5000, $\rho$ = -0.5 obs = 10000, $\rho$ = -0.5
$\beta_{0}$ $2,860.340$ $310.396$ $61.885$
$\beta_{1}$ $8,978.860$ $2,086.819$ $816.737$
$\pi_{0}$ $2,553.201$ $688.355$ $2.279$
$\pi_{1}$ $7,481.836$ $2,942.766$ $4.219$
$\lambda$ $7,083.483$ $1,754.262$ $714.109$
$\zeta$ $5,732.446$ $2,576.872$ $2.182$
$\rho$ $0.650$ $0.575$ $0.489$
Case 1-a
Param. obs = 2000, $\rho$ = -0.5 obs = 5000, $\rho$ = -0.5 obs = 10000, $\rho$ = -0.5
$\beta_{0}$ $684.210$ $0.190$ $0.126$
$\beta_{1}$ $3,169.837$ $0.471$ $0.285$
$\pi_{0}$ $2,328.631$ $1.059$ $0.736$
$\pi_{1}$ $1,702.444$ $0.685$ $0.474$
$\pi_{2}$ $1,661.837$ $0.433$ $0.293$
$\lambda$ $2,992.510$ $0.390$ $0.239$
$\zeta$ $1,925.182$ $0.607$ $0.428$
$\rho$ $0.488$ $0.362$ $0.233$
Case 1-b
Param. obs = 2000, $\rho$ = -0.5 obs = 5000, $\rho$ = -0.5 obs = 10000, $\rho$ = -0.5
$\beta_{0}$ $5,103.763$ $211.961$ $0.187$
$\beta_{1}$ $11,215.290$ $608.039$ $0.436$
$\pi_{0}$ $3,290.975$ $1.583$ $1.146$
$\pi_{1}$ $2,035.018$ $0.756$ $0.561$
$\pi_{2}$ $1,936.292$ $0.576$ $0.341$
$\lambda$ $13,803.720$ $833.790$ $0.522$
$\zeta$ $3,457.443$ $1.040$ $0.719$
rho $0.595$ $0.505$ $0.400$
Case 2-a
Param. obs = 2000, $\rho$ = -0.5 obs = 5000, $\rho$ = -0.5 obs = 10000, $\rho$ = -0.5
$\beta_{0}$ $0.323$ $0.165$ $0.110$
$\beta_{1}$ $0.621$ $0.265$ $0.175$
$\beta_{2}$ $0.549$ $0.233$ $0.150$
$\pi_{0}$ $616.635$ $0.476$ $0.335$
$\pi_{1}$ $438.369$ $0.359$ $0.249$
$\pi_{2}$ $484.978$ $0.308$ $0.215$
$\lambda$ $0.712$ $0.291$ $0.191$
$\zeta$ $587.512$ $0.415$ $0.288$
$\rho$ $0.171$ $0.103$ $0.072$
Case 2-b
Param. obs = 2000, $\rho$ = -0.5 obs = 5000, $\rho$ = -0.5 obs = 10000, $\rho$ = -0.5
$\beta_{0}$ $2,581.494$ $107.814$ $0.164$
$\beta_{1}$ $4,770.145$ $256.687$ $0.261$
$\beta_{2}$ $4,687.187$ $222.714$ $0.223$
$\pi_{0}$ $3,211.289$ $0.958$ $0.587$
$\pi_{1}$ $1,913.604$ $0.512$ $0.305$
$\pi_{2}$ $2,048.146$ $0.454$ $0.261$
$\lambda$ $8,605.971$ $459.471$ $0.410$
$\zeta$ $3,794.507$ $0.889$ $0.509$
$\rho$ $0.255$ $0.146$ $0.104$
Notes: See Table 1 for the specific design of each case. All results are produced using $2000$ Monte Carlo replications.
Analyzing the results presented in Table 2, we observe a pronounced MB across all scenarios for sample sizes of $n=2000$ and $n=5000$, which substantially diminishes at $n=10000$. This behavior aligns with the findings for the RMSE in Table 3. In Case 0, the absence of exclusion restrictions results in a persistent identification issue across all sample sizes, as evidenced by the excessive MB and RMSE. Notably, scenarios with a single (continuous) excluded variable in only one of the equations, specifically Cases 1-a and 1-b, also demonstrate substantial MB at $n=10000$, especially for intercepts, scaling parameters, and dependence parameters. Conversely, in Cases 2-a and 2-b, which feature two exclusion restrictions (with continuous excluded variables), MB and RMSE are substantially reduced compared to their counterparts in Cases 1-a and 1-b. Interestingly, the (unconditional) variance of the index emerges as a crucial factor, as demonstrated by the larger MB and RMSE in Cases 1-b and 2-b, where the index variance is substantially smaller than that of the error. [^12] The RMSE results for Cases 1-b and 2-b indicate that various parameters are very imprecisely estimated at smaller sample sizes, often attributed to outlier estimates. A detailed investigation reveals that this imprecision often originates from outlier estimates, further emphasizing the importance of a large index variance for precise estimation.
In summary, our findings from this section suggest that employing at least two exclusion restrictions (one in each equation) effectively reduces the finite sample MB and RMSE when estimating model parameters of the joint distribution in a partially observable framework. Furthermore, a larger index variance contributes to precise estimation.
@@ 9 Empirical Application
On the contrary, various other countries have initiated a legalization process of the possession and consumption of certain categories of cannabis drugs for recreational and medicinal purposes, driven by aims to reduce drug-related crime and corresponding law enforcement costs, as well as to generate additional tax revenue. Notably, nearly one-quarter of the U.S. states have legalized recreational marijuana by now, regulating and taxing its sales (Brinkman and Mok-Lamme, 2019; Hansen et al., 2020).
@@
The outcome variable of interest (cons_30days), based on the survey responses to the question “On how many occasions have you used marijuana (weed, pot) or hashish (hash, hash oil) … in the last 30 days?”, is categorized into three levels: “0” (never used), “1” (used 1-2 times), and “2” (used more than 2 times). Similar to Hollingworth et al. (2022), we not only include individual-level covariates like gender, age, ethnicity, and parental education, but also certain state-level control variables such as unemployment rate and median income, fixed at pre-treatment levels. [^16] State-level unemployment rates as well as the state median income were constructed from data of the Bureau of Labor Statistics (https://www.bls.gov/). Finally, we employ state-level marijuana prices from 2015 as an instrument for the consumption (state_price_m), and a “survey cooperation” measure (trust) for the reporting, following Greene et al. (2018) and Brown et al. (2018). [^17] Marijuana price data source:
@@
Table 4 presents descriptive statistics of the outcome variable and other controls. As expected, marijuana consumption among 8th graders appears relatively low, with only 6% of the sample reporting usage over the past 30 days. Further, unreported results indicate that these characteristics remain largely stable across time within each group.
Table 4: Summary Statistics
Variable N Mean Std. Dev. Min Pctl. 25 Pctl. 75 Max
cons_30days 46472
… 1-2 occasions 1290 3%
… 3 or more occasions 1375 3%
… no use 43807 94%
male 46472 0.48 0.5 0 0 1 1
age_month 46472 169 6.2 122 165 173 231
white 46472 0.47 0.5 0 0 1 1
parent_edu_clg 46472 0.58 0.49 0 0 1 1
state_unemp 46472 5.3 0.78 2.9 4.8 5.9 6.8
state_median_income 46472 56552 7024 40037 51983 60925 75675
state_price_m 46472 240 35 159 220 265 325
trust 46472 0.96 0.072 0.28 0.96 1 1
treated 46472 0.19 0.39 0 0 0 1
after 46472 0.47 0.5 0 0 1 1
Notes: Data for 8th graders from 2015-2018, excluding the states Oregon, District of Columbia, Colorado and Washington. treated = 1 if the respondent is from California, Maine, Massachusetts and Nevada, = 0 otherwise. after = 1 if the respondent is surveyed in 2017-2018, = 0 otherwise. The variables male, white, parent_edu_clg are binary indicators for self-reported gender status (male), ethnicity (white), and parental education level (college) equal to one if the category was true, and zero otherwise. For the other variables, please refer to the text for the definition.
In Table 5, we present our initial findings from the unconditional analysis, as defined in (21), without misreporting and covariates. Our first set of results (gdid) suggests that $\tau^{c}_{11}(j)$ is not statistically different from zero for $j\in\{0,1,2\}$ at any conventional levels. These results align with the estimates from the more restrictive ordered (pdid_ordered) and binary choice (pdid_probit) DiD models, which include the binary time period and group indicators and their interaction as regressors.
@@
Data between 2015-2018 are used.  The first panel of results (“Without Misreporting”) displays estimates of treatment effects $\tau_{11}(j)$, $j\in\{0,1,2\}$, for the model estimated without misreporting. The treatment effect estimates have been marginalized with respect to the distribution of $\mathbf{X}_{11}$ (see Equation (21)). The second panel of results (“With Misreporting”) displays estimates of treatment effects $\tau_{11}(j)$ and $\tau_{11}^{r}(j)$, $j\in\{0,1,2\}$, for the model with misreporting. The treatment effect estimates have also been marginalized with respect to the distribution of $\mathbf{X}_{11}$. The first two columns display results for a Frank Copula specification, the last two columns show corresponding results for a Clayton Copula specification. Standard errors are computed via the Delta method and can be found in the corresponding columns next to the coefficient estimates.

### v3

@@ 4.2.2 Distributional Treatment Effects
$H_{A}:\overline{\boldsymbol{\eta}}^{(0)}_{1t}-\overline{\boldsymbol{\eta}}^{(0)}_{0t}-\frac{\lambda^{(0)}_{0t}}{\lambda^{(0)}_{0(t-1)}}\left(\overline{\boldsymbol{\eta}}^{(0)}_{1(t-1)}-\overline{\boldsymbol{\eta}}^{(0)}_{0(t-1)}\right)=\mathbf{0}\quad\text{and/ or}\quad\frac{\lambda^{(0)}_{1t}}{\lambda^{(0)}_{1(t-1)}}-\frac{\lambda^{(0)}_{0t}}{\lambda^{(0)}_{0(t-1)}}=0.$
via a Likelihood Ratio test comparing an unrestricted log-likelihood as in Equation (15) with a restricted version that satisfies $\overline{\boldsymbol{\eta}}^{(0)}_{1t}=\overline{\boldsymbol{\eta}}^{(0)}_{0t}+\frac{\lambda^{(0)}_{0t}}{\lambda^{(0)}_{0(t-1)}}(\overline{\boldsymbol{\eta}}^{(0)}_{1(t-1)}-\overline{\boldsymbol{\eta}}^{(0)}_{0(t-1)})$ and $\lambda^{(0)}_{1t}=\frac{\lambda^{(0)}_{0t}\lambda^{(0)}_{1(t-1)}}{\lambda^{(0)}_{0(t-1)}}$, respectively.
@@ 5 Empirical Application
The outcome variable of interest (cons_30days), based on the survey responses to the question “On how many occasions have you used marijuana (weed, pot) or hashish (hash, hash oil) … in the last 30 days?”, is categorized into three levels: “0” (never used), “1” (used 1-2 times), and “2” (used more than 2 times).  [^11] Table LABEL:Summary1 in Section LABEL:sec:additionalempirical of the supplement presents descriptive statistics of the outcome variable and other controls. Similar to Hollingworth et al. (2022), we not only include individual-level covariates like gender, age, ethnicity, and parental education, but also certain state-level control variables such as unemployment rate and median income, fixed at pre-treatment levels. [^12] State-level unemployment rates as well as the state median income were constructed from data of the Bureau of Labor Statistics (https://www.bls.gov/). Finally, we employ state-level marijuana prices from 2015 as an instrument for the consumption (state_price_m), and a “survey cooperation” measure (trust) for the reporting, following Greene et al. (2018) and Brown et al. (2018). [^13] Marijuana price data source:
@@
In Table 1, we present our initial findings for a model without underreporting and without additional control variables. Our first set of results (gdid) suggests that $\tau^{c}_{11}(j)$ is not statistically different from zero for $j\in\{0,1,2\}$ at any conventional levels. These results align with the estimates from the more restrictive ordered (pdid_ordered) and binary choice (pdid_probit) DiD models, which include the binary time period and group indicators and their interaction as regressors.
@@
Data between 2015-2018 are used.  The first panel of results (“Without Misreporting”) displays estimates of treatment effects $\tau_{11}(j)$, $j\in\{0,1,2\}$, for the model estimated without misreporting. The treatment effect estimates have been marginalized with respect to the distribution of $\mathbf{X}_{11}$. The second panel of results (“With Misreporting”) displays estimates of treatment effects $\tau_{11}(j)$ and $\tau_{11}^{r}(j)$, $j\in\{0,1,2\}$, for the model with misreporting. The treatment effect estimates have also been marginalized with respect to the distribution of $\mathbf{X}_{11}$. The first two columns display results for a Frank Copula specification, the last two columns show corresponding results for a Clayton Copula specification. Standard errors are computed via the Delta method and can be found in the corresponding columns next to the coefficient estimates.